# UAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers

### v1

@@ D.1 Sensitivity Analysis of and
We examine the sensitivity of UAT-Lite to its two primary inference-time hyperparameters: the uncertainty weighting coefficient and the number of Monte Carlo samples . A small, representative grid is evaluated with and on SQuAD 2.0, which is particularly sensitive to epistemic uncertainty due to the presence of both answerable and unanswerable questions. Similar stability trends were observed on MNLI and SST-2 and are omitted for brevity. We report ECE and predictive accuracy to assess both calibration quality and task performance. Accuracy is reported here solely to verify that calibration gains do not arise from performance degradation.
Across all configurations, calibration performance remains stable, with ECE values confined to a narrow interval between 0.064 and 0.070 (mean , standard deviation ). No sharp sensitivity to either or is observed. Increasing the number of Monte Carlo samples yields only marginal calibration improvements beyond , indicating diminishing returns from additional stochastic passes.
Predictive accuracy remains effectively unchanged across the grid, varying by less than 0.5% absolute, suggesting that uncertainty-weighted attention primarily affects calibration rather than accuracy. Overall, these results indicate that UAT-Lite does not rely on precise hyperparameter tuning to achieve effective calibration and is robust across a broad range of reasonable inference-time settings. This stability supports the practicality of the method in deployment scenarios where extensive hyperparameter search is undesirable.
| ECE | Accuracy | ||
| 0.1 | 3 | 0.0699 | 0.7075 |
| 0.1 | 5 | 0.0676 | 0.7075 |
| 0.1 | 10 | 0.0649 | 0.7109 |
| 0.5 | 3 | 0.0703 | 0.7059 |
| 0.5 | 5 | 0.0657 | 0.7096 |
| 0.5 | 10 | 0.0682 | 0.7068 |
| 1.0 | 3 | 0.0669 | 0.7090 |
| 1.0 | 5 | 0.0655 | 0.7102 |
| 1.0 | 10 | 0.0640 | 0.7110 |
Summary. ECE mean , std ; min , max (range ).
D.2 Clinical Domain Calibration
Table 7 reports calibration and accuracy on two clinical benchmarks with distinct task structures and uncertainty profiles: MedQA and PubMedQA. We emphasize that these clinical datasets are used only to assess domain transfer of uncertainty-aware attention to medically specialized language distributions, and should not be interpreted as evidence of clinical decision reliability. In particular, what constitutes “good uncertainty” in clinical settings is often decision-dependent and may involve asymmetric error costs (e.g., prioritizing recall for critical conditions), which are not explicitly modeled in our evaluation.
On MedQA, Deep Ensembles achieve the lowest ECE (0.007) and the highest accuracy (0.268), reflecting the benefit of model diversity in discrete, fact-centric multiple-choice medical question answering. In contrast, UAT-Lite exhibits a slightly higher ECE (0.040) with marginal accuracy improvement over the baseline. This behavior is expected, as MedQA primarily evaluates factual recall with limited contextual ambiguity, constraining the benefits of uncertainty-aware attention. Importantly, UAT-Lite achieves competitive performance without the substantial computational and storage overhead associated with training and maintaining multiple model replicas.
A different trend emerges on PubMedQA. While Deep Ensembles substantially reduce ECE (0.051), this improvement coincides with a marked drop in accuracy (0.520), indicating overly conservative confidence estimates. In contrast, UAT-Lite attains the highest accuracy (0.640) while maintaining calibration comparable to the baseline. This suggests that integrating uncertainty directly into the attention mechanism can improve evidence aggregation in long, semantically dense biomedical abstracts, where uncertainty arises from contextual interpretation rather than discrete fact retrieval.
Overall, these results highlight a trade-off between ensemble-based calibration and inference-time uncertainty integration. While ensembles excel in fact-centric clinical QA, UAT-Lite offers a more favorable accuracy–calibration balance on tasks characterized by semantic ambiguity and long-context reasoning, without incurring additional training or parameter costs. Accordingly, we treat the clinical-domain results as a domain-transfer generalization stress test rather than a primary claim about clinical deployment.
| Dataset | Method | ECE | Accuracy |
| MedQA | Baseline | 0.033 | 0.246 |
| Temperature Scaling | 0.023 | 0.246 | |
| Deep Ensemble (5) | 0.007 | 0.268 | |
| UAT-Lite | 0.040 | 0.249 | |
| PubMedQA | Baseline | 0.114 | 0.620 |
| Temperature Scaling | 0.128 | 0.620 | |
| Deep Ensemble (5) | 0.051 | 0.520 | |
| UAT-Lite | 0.103 | 0.640 |
D.3 Character-Level Adversarial Robustness
Table 8 evaluates robustness under character-level perturbations on SST-2. We apply random 5% character swaps to simulate typographical errors and adversarial perturbations.
@@ D.3 Character-Level Adversarial Robustness
Both baseline BERT and UAT-Lite exhibit substantial accuracy degradation under attack, with baseline accuracy dropping from 0.920 to 0.816 and UAT-Lite from 0.922 to 0.812. Calibration also degrades substantially, with ECE increasing from 0.062 to 0.133 for baseline and from 0.055 to 0.132 for UAT-Lite.
While UAT-Lite maintains slightly better clean accuracy (0.922 vs. 0.920) and clean calibration (ECE 0.055 vs. 0.062), the robustness under attack is comparable between methods. This indicates that uncertainty-aware attention improves calibration on clean data but provides limited additional protection against adversarial perturbations.
These results are consistent with prior work showing that calibration methods address epistemic uncertainty but do not directly defend against adversarial attacks Goodfellow et al. (2014). Future work should explore combining UAT-Lite with specialized adversarial defenses such as adversarial training or certified robustness techniques.
@@
D.4 Full Linguistic Uncertainty Analysis
Table 9 provides a fine-grained linguistic analysis of how UAT-Lite responds to diverse sources of uncertainty, spanning lexical, syntactic, epistemic, and logical phenomena.
For negation, UAT-Lite remains uniformly confident across affirmative, negated, and clinical negation cases, yielding zero estimated uncertainty. This behavior indicates negation invariance, which is desirable for explicit clinical negation (e.g., “not stable”) but also suggests limited sensitivity to polarity reversal in non-clinical contexts. This invariance is consistent with the attention analysis in Fig.˜3, which shows that uncertainty-aware attention redistributes weight conservatively toward negation and predicate structure without inducing elevated output-level uncertainty for clear clinical negation.
In cases of lexical ambiguity, the framework exhibits mixed behavior. When sufficient contextual disambiguation is present (“bank account”), UAT-Lite appropriately increases uncertainty. However, for genuinely ambiguous expressions (“bank closed”), the model remains highly confident, indicating overconfidence under unresolved polysemy. Partial sensitivity is observed for polysemous terms such as “bat,” where uncertainty increases but does not fully reflect ambiguity.
A similar pattern arises for epistemic hedges. Strong assertions (“definitely”) and weak hedges (“probably,” “possibly”) are treated with near-maximal confidence, suggesting hedge insensitivity. In contrast, explicit modal uncertainty (“might”) leads to a measurable increase in uncertainty, indicating that the model responds primarily to strong lexical uncertainty markers rather than gradated epistemic cues.
For syntactic complexity, UAT-Lite demonstrates heightened sensitivity. Medium-depth syntactic structures produce the largest uncertainty increase, reflecting increased structural processing difficulty. Interestingly, deeper structures elicit a non-monotonic response, suggesting that the model may partially adapt to highly regular but complex constructions rather than uniformly increasing uncertainty with depth.
Finally, under logical contradictions, UAT-Lite appropriately increases uncertainty for inconsistent inputs (“healthy dying”) while maintaining high confidence for coherent statements. This indicates effective detection of semantic inconsistency, consistent with the framework’s uncertainty propagation across layers.
@@ Appendix E Cross-Model & Scaling Analysis
Scope clarification. All models evaluated in this section are encoder-only, BERT-derived transformer architectures. Generalization to fundamentally different architectures (e.g., decoder-only models such as GPT or LLaMA, or state-space models) is outside the scope of the current study and is left for future work.
@@ E.1 Generalization Across BERT Backbones
Table 10 evaluates whether the calibration benefits of UAT-Lite generalize across BERT-based transformer backbones pretrained on heterogeneous corpora and applied to domain-specific benchmarks. The evaluated models span general-purpose (BERT-base, RoBERTa-base), clinical (ClinicalBERT), biomedical (BioBERT), and scientific (SciBERT) representations, thereby testing robustness across substantially different linguistic and semantic distributions.
Across evaluated domains, UAT-Lite reduces ECE for most pretrained backbones, yielding an average relative improvement of approximately 32% (Table 10).
The strongest gain is observed for BERT-base on SQuAD, where ECE decreases by 52% (0.130 0.062). This result is particularly significant because SQuAD contains both answerable and unanswerable questions, making it highly sensitive to epistemic uncertainty. The large improvement indicates that uncertainty-weighted attention enables the model to better recognize ambiguity during evidence aggregation rather than producing overconfident predictions.
In domain-specialized settings, BioBERT on PubMedQA achieves a 40% reduction in ECE, suggesting that uncertainty-aware attention is especially beneficial for long, evidence-dense biomedical text that requires integrating information across multiple sentences. ClinicalBERT on MedQA shows a 25% improvement, indicating that the method remains effective even for fact-oriented medical question answering, although gains are smaller due to relatively strong baseline calibration.
By contrast, RoBERTa-base on MNLI exhibits only a marginal 2% improvement. MNLI is a large, high-resource dataset with strong pretrained representations and comparatively well-calibrated baselines, leaving limited room for additional calibration gains. SciBERT on SQuAD similarly shows a modest 10% reduction, suggesting that pretraining on highly specialized scientific corpora may already encode task-relevant uncertainty patterns.
@@ E.2 Effect of Model Capacity on Calibration
Table 11 analyzes the effect of model capacity on calibration improvement by applying UAT-Lite across different sizes of the BERT architecture, ranging from BERT-tiny (4.4M parameters) to BERT-large (335.1M parameters). Unlike Table 10, which varies domains and pretraining corpora, this analysis holds the architecture family fixed and isolates the role of parameter scale.
Small and mid-sized models benefit substantially from uncertainty-weighted attention. BERT-mini achieves the largest relative improvement, with a 58% reduction in ECE, followed by BERT-base with an approximately 53% improvement (52–53% depending on rounding and binning resolution). These models exhibit sufficient representational capacity to express epistemic uncertainty, while still relying on attention mechanisms that benefit from explicit uncertainty modulation.
BERT-tiny and BERT-small show more modest improvements (12% and 15%, respectively). In very small models, limited representational capacity constrains the extent to which uncertainty-aware reasoning can reshape internal attention patterns, resulting in smaller but still consistent calibration gains.
In contrast, BERT-large exhibits degraded calibration, with ECE increasing by 13% when UAT-Lite is applied. We hypothesize that in highly overparameterized models, stochastic perturbations introduced via Monte Carlo sampling may disrupt already well-formed attention patterns, leading to excessive variance propagation rather than meaningful uncertainty modeling. This observation suggests that uncertainty-weighted attention is most effective in the regime of small to mid-sized transformers, where model capacity and stochasticity appear to be better balanced.
Latency measurements further contextualize this behavior. While inference-time overhead remains moderate for mid-sized models (e.g., 7 ms 39 ms for BERT-base), it increases substantially for larger architectures (26 ms 134 ms for BERT-large), reinforcing the trade-off between calibration gains and computational cost.
@@ Appendix F Implementation and Reproducibility
This section characterizes the computational and implementation properties of UAT-Lite. We first analyze its asymptotic and practical computational complexity relative to standard transformer inference and ensemble-based uncertainty methods. We then quantify inference-time overhead introduced by Monte Carlo sampling and describe the experimental setup, hardware configuration, and reproducibility details required to replicate all reported results. Throughout, we emphasize that UAT-Lite operates strictly at inference time and does not modify pretrained parameters or training cost.
@@ F.1 Computational Complexity Analysis
UAT-Lite preserves the parameter count and training cost of the base transformer. All uncertainty-aware mechanisms operate strictly at inference time and do not introduce additional trainable parameters or modify learned weights.
Additional computational cost arises solely from Monte Carlo (MC) sampling during inference:
@@
Inference cost: stochastic forward passes (typically )
From a complexity perspective, inference-time cost scales linearly with the number of Monte Carlo samples . In contrast to ensemble-based approaches, which require training, storing, and evaluating multiple independent models, UAT-Lite incurs no additional training or memory overhead. As a result, its total computational footprint remains substantially lower than Deep Ensembles for comparable calibration gains.
@@ F.2 Inference-Time Overhead
While the previous subsection characterizes asymptotic complexity, we now report empirical inference-time overhead under realistic deployment conditions.
At test time, UAT-Lite performs stochastic forward passes through the same pretrained model. When executed sequentially, runtime increases approximately linearly with . For BERT-base, this corresponds to a measured slowdown of approximately relative to deterministic inference when using .
Despite this overhead, inference-time cost remains practical for many deployment settings, particularly in scenarios where reliability and calibrated confidence are prioritized over raw throughput. Importantly, this cost is incurred only when uncertainty-aware inference is enabled and can be adjusted dynamically by reducing for high-confidence inputs, as discussed in Section Limitations.
@@ F.3 Reproducible Experimental Setup
Learning rate schedule: linear warmup for first 500 steps, followed by linear decay
@@
Inference via Monte Carlo sampling with stochastic forward passes
@@
Sensitivity analysis for and reported in §˜D.1
@@ F.4 Hardware, Latency, and Code Availability
All experiments are conducted on a single NVIDIA A100 GPU (40GB). Unless otherwise stated, latency is measured as end-to-end forward-pass wall-clock time per example, averaged over 200 runs after 50 warm-up iterations. We use fixed sequence lengths (128 for SST-2 and MNLI; task-standard truncation for SQuAD answerability) and batch size 1.
For stochastic methods, latency is reported under the same Monte Carlo budget used during evaluation. Reported slowdowns reflect the ratio between stochastic inference and deterministic baseline inference under identical conditions.
@@
To ensure reproducibility during the anonymous review process, we provide an anonymized repository containing the complete implementation, including training scripts, evaluation pipelines, and configuration files. All experiments are implemented using HuggingFace Transformers, PyTorch, and Python. The repository is available at: https://github.com/eliashossain001/uq_decomposition/tree/main.

### v2

@@ D.11 Sensitivity Analysis of and
We examine the sensitivity of UAT-Lite to two inference-time hyperparameters: the uncertainty weighting coefficient and the Monte Carlo budget . We evaluate a representative grid with and on SQuAD 2.0, which is particularly sensitive to epistemic uncertainty due to the presence of both answerable and unanswerable questions. Similar stability trends were observed on MNLI and SST-2 and are omitted for brevity. We report ECE and accuracy to assess calibration quality and task performance; accuracy is included to verify that calibration changes are not driven by degraded performance.
Table 13 shows that calibration remains stable across the grid: ECE lies in a narrow interval (mean , std ). No sharp sensitivity to is observed; across blocks, ECE varies by at most . Within each , increasing tends to reduce ECE relative to the reference (ECE is consistently non-positive), but gains diminish beyond : moving from yields most of the improvement at a moderate cost increase (MC Cost ), while provides smaller additional benefits despite a cost. Accuracy is similarly robust: changes within each block are small (Acc absolute), indicating that uncertainty-weighted attention primarily modulates calibration rather than task accuracy. Overall, the method is insensitive to precise tuning of and achieves stable calibration over a broad range of inference-time settings, supporting practicality when extensive hyperparameter search is undesirable.
ECE Acc ECE Acc MC Cost (vs. ) (vs. ) (vs. ) 0.1 3 0.0699 0.7075 0.0000 0.0000 1.00 0.1 5 0.0676 0.7075 1.67 0.1 10 0.0649 0.7109 3.33 0.5 3 0.0703 0.7059 0.0000 0.0000 1.00 0.5 5 0.0657 0.7096 1.67 0.5 10 0.0682 0.7068 3.33 1.0 3 0.0669 0.7090 0.0000 0.0000 1.00 1.0 5 0.0655 0.7102 1.67 1.0 10 0.0640 0.7110 3.33
Summary. ECE mean , std ; min , max (range ). ECE/Acc are computed within each block relative to ; MC Cost scales linearly with .
D.12 Character-Level Adversarial Robustness
@@
Table 14 evaluates robustness under character-level perturbations on SST-2. We apply random 5% character swaps to simulate typographical noise and simple adversarial corruption. Both the baseline BERT model and UAT-Lite exhibit substantial accuracy degradation under attack, with accuracy dropping from 0.920 to 0.816 for the baseline and from 0.922 to 0.812 for UAT-Lite. Calibration also deteriorates markedly, with ECE increasing from 0.062 to 0.133 for the baseline and from 0.055 to 0.132 for UAT-Lite.
Although UAT-Lite achieves slightly better clean accuracy (0.922 vs. 0.920) and clean calibration (ECE 0.055 vs. 0.062), its robustness under attack remains broadly comparable to that of the baseline. These results suggest that uncertainty-aware attention improves calibration under clean conditions but provides limited additional robustness to character-level perturbations. This pattern is consistent with prior work showing that calibration-oriented methods primarily address uncertainty estimation and do not, by themselves, constitute adversarial defenses Goodfellow et al. (2014). An important direction for future work is to combine UAT-Lite with dedicated adversarial-defense strategies, such as adversarial training or certified robustness methods.
@@
D.13 Full Linguistic Uncertainty Analysis
Table 15 provides a fine-grained analysis of how UAT-Lite responds to diverse sources of uncertainty spanning lexical, syntactic, epistemic, and logical phenomena. For negation, UAT-Lite remains uniformly confident across affirmative, negated, and clinical-negation cases, yielding near-zero estimated uncertainty. This behavior indicates invariance to explicit negation cues, which is desirable for clear clinical negation (e.g., “not stable”) but also suggests limited sensitivity to polarity reversal in more general contexts. This pattern is consistent with the attention analysis in Fig.˜7, which shows that uncertainty-aware attention redistributes weight conservatively toward negation and predicate structure without producing elevated output-level uncertainty for clear clinical negation.
For lexical ambiguity, the behavior is mixed. When contextual disambiguation is available (“bank account”), UAT-Lite shows a meaningful increase in uncertainty. However, for genuinely ambiguous expressions (“bank closed”), the model remains highly confident, indicating residual overconfidence under unresolved polysemy. A similar intermediate pattern appears for polysemous triggers such as “bat,” where uncertainty increases but does not fully reflect the underlying ambiguity.
A related pattern appears for epistemic hedges. Strong assertions (“definitely”) and weaker hedges (“probably,” “possibly”) are treated with near-maximal confidence, indicating limited sensitivity to graded epistemic marking. By contrast, explicit modal uncertainty (“might”) leads to a measurable increase in uncertainty, suggesting that the model responds more strongly to direct lexical uncertainty markers than to subtler hedging cues.
For syntactic complexity, UAT-Lite shows clearer sensitivity. Medium-depth structures produce the largest increase in uncertainty, consistent with greater structural processing difficulty. Deeper structures, however, elicit a non-monotonic response, suggesting that the model may partially adapt to highly regular but more complex constructions rather than increasing uncertainty monotonically with depth. Finally, under logical contradictions, UAT-Lite appropriately increases uncertainty for inconsistent inputs (“healthy dying”) while maintaining high confidence for coherent statements. This pattern indicates sensitivity to semantic inconsistency and is consistent with the framework’s uncertainty propagation across layers.
@@ Appendix E Cross-Model & Scaling Analysis
Scope clarification. All models evaluated in this section are encoder-only, BERT-derived transformer architectures. Generalization to fundamentally different architectures, such as decoder-only models (e.g., GPT or LLaMA) or state-space models, is outside the scope of the current study and is left for future work.
@@ E.1 Generalization Across BERT Backbones
Table 16 evaluates whether the calibration benefits of UAT-Lite generalize across BERT-based transformer backbones pretrained on heterogeneous corpora and applied to domain-specific benchmarks. The evaluated models span general-purpose (BERT-base, RoBERTa-base), clinical (ClinicalBERT), biomedical (BioBERT), and scientific (SciBERT) representations, thereby testing robustness across substantially different linguistic and semantic domains.
Across the evaluated settings, UAT-Lite reduces ECE for most pretrained backbones, yielding an average relative improvement of approximately 32% (Table 16). The largest gain is observed for BERT-base on SQuAD, where ECE decreases by 52% (0.130 0.062). Because SQuAD contains both answerable and unanswerable questions, it places substantial demands on ambiguity recognition and confidence estimation. The large improvement in this setting suggests that uncertainty-weighted attention helps the model better calibrate predictions when evidence aggregation is uncertain.
Domain-specialized models also benefit, although to different degrees. BioBERT on PubMedQA shows a 40% reduction in ECE, suggesting that uncertainty-aware attention is useful for evidence-dense biomedical text. ClinicalBERT on MedQA shows a 25% improvement, indicating that the method remains effective in fact-oriented medical question answering, although the gain is smaller than in some other settings. By contrast, RoBERTa-base on MNLI shows only a marginal 2% improvement. MNLI is a high-resource benchmark with strong pretrained representations and a relatively well-calibrated baseline, leaving less room for further gains. SciBERT on SQuAD similarly shows a modest 10% reduction, suggesting that pretraining on specialized scientific corpora may already capture some task-relevant uncertainty structure.
Taken together, these results indicate that the calibration benefits of UAT-Lite transfer across multiple BERT-derived backbones and application domains, with the largest gains arising in settings characterized by greater ambiguity or less favorable baseline calibration.
@@ E.2 Effect of Model Capacity on Calibration
Table 17 analyzes the effect of model capacity on calibration improvement by applying UAT-Lite across BERT models ranging from BERT-tiny (4.4M parameters) to BERT-large (335.1M parameters). Unlike Table 16, which varies domains and pretraining corpora, this analysis holds the architecture family fixed and isolates the role of parameter scale.
Small and mid-sized models benefit most from uncertainty-weighted attention. BERT-mini achieves the largest relative improvement, with a 58% reduction in ECE, followed by BERT-base with an approximately 53% reduction. These models appear to have sufficient representational capacity to express epistemic uncertainty while still benefiting from explicit uncertainty modulation within attention.
BERT-tiny and BERT-small show more modest improvements (12% and 15%, respectively). For very small models, limited representational capacity may constrain the extent to which uncertainty-aware inference can reshape internal attention patterns, yielding smaller but still consistent gains. In contrast, BERT-large shows degraded calibration, with ECE increasing by 13% after applying UAT-Lite. One possible explanation is that, in highly overparameterized models, stochastic perturbations introduced through Monte Carlo sampling may disrupt already stable attention patterns, leading to excess variance propagation rather than improved uncertainty estimation.
Latency measurements further contextualize these results. Although inference-time overhead remains moderate for mid-sized models (e.g., 7 ms 39 ms for BERT-base), it increases substantially for larger architectures (26 ms 134 ms for BERT-large). Overall, this analysis suggests that uncertainty-weighted attention is most effective in the small-to-mid-sized regime, where model capacity and stochasticity appear to be better balanced.
@@ Appendix F Implementation and Reproducibility
This section characterizes the computational and implementation properties of UAT-Lite. We first analyze its asymptotic and practical computational complexity relative to standard transformer inference and ensemble-based uncertainty methods. We then quantify the inference-time overhead introduced by Monte Carlo sampling and summarize the experimental setup, hardware configuration, and reproducibility details required to replicate the reported results. Throughout, we emphasize that UAT-Lite operates strictly at inference time and does not alter pretrained parameters or training cost.
@@ F.1 Computational Complexity Analysis
UAT-Lite preserves the parameter count and training cost of the base transformer. All uncertainty-aware mechanisms operate strictly at inference time and introduce no additional trainable parameters or architectural modifications.
The additional cost arises solely from Monte Carlo (MC) sampling during inference:
@@
Inference cost: stochastic forward passes ()
Inference-time complexity therefore scales linearly with the MC budget . Table 18 reports a canonical end-to-end GPU benchmark at under a sequential MC evaluation loop. Relative to deterministic inference, UAT-Lite incurs approximately a slowdown, whereas temperature scaling adds negligible overhead because it applies only a scalar rescaling to the logits. Unlike Deep Ensembles, which require training, storing, and evaluating multiple independent models, UAT-Lite introduces no additional training or model-storage cost, making it a lighter-weight alternative when retraining is undesirable.
| Method | Latency (ms) | Ratio | |
| Base (det.) | 62.93 | 0.29 | 1.00 |
| Base + TS | 63.13 | 0.27 | 1.00 |
| UAT-Lite (M=10) | 1426.90 | 78.87 | 22.68 |
| UAT-Lite + TS (M=10) | 1486.61 | 7.26 | 23.62 |
@@ F.2 Inference-Time Overhead
While the previous subsection characterizes asymptotic complexity, we now report empirical inference-time overhead under realistic deployment conditions. At test time, UAT-Lite performs stochastic forward passes through the same pretrained model. When executed sequentially, runtime increases approximately linearly with . Observed slowdowns depend on the Monte Carlo budget and the latency measurement protocol, including batching, sequence length, and whether MC passes are executed sequentially or vectorized. Throughout this work, we report latency under the canonical benchmark described in F.4, under which overhead scales approximately linearly with for sequential MC inference.
Despite this overhead, inference-time cost remains practical in settings where reliability and calibrated confidence are prioritized over raw throughput. Importantly, this cost is incurred only when uncertainty-aware inference is enabled and can in principle be reduced by lowering in settings where a smaller stochastic budget is acceptable, as discussed in Section Limitations.
@@ F.3 Reproducible Experimental Setup
Learning rate schedule: linear warmup for the first 500 steps, followed by linear decay
@@
Inference via Monte Carlo sampling with stochastic forward passes (default for the main revision tables; is additionally reported where noted)
@@
Sensitivity analysis for and reported in D.11
@@ F.4 Hardware, Latency, and Code Availability
All experiments are conducted on a single NVIDIA A100 GPU (40GB). Unless otherwise stated, latency is measured as end-to-end forward-pass wall-clock time, averaged over 200 runs after 50 warm-up iterations. We report a single canonical latency benchmark throughout the paper: batch size , sequence length , a sequential MC loop without vectorization across MC samples, and a single GPU. Latency is reported per batch and expressed as a ratio relative to deterministic inference under the same conditions. For stochastic methods, latency is measured using the same Monte Carlo budget used during evaluation. Reported slowdowns therefore reflect the ratio between stochastic and deterministic inference under identical measurement settings.
@@
To support reproducibility during anonymous review, we provide an anonymized repository containing the complete implementation, including training scripts, evaluation pipelines, and configuration files. All experiments are implemented using HuggingFace Transformers, PyTorch, and Python. The repository is available at: https://github.com/eliashossain001/uq_decomposition/tree/main.